Please help me solve this;46% OF ADULTS SAY CASHEWS ARE THEIR FAVORITE KIND OF NUT. YOU RANDOMLY SELECT 12 ADULTS AND ASK EACH TO NAME HIS OR HER FAVORITE NUT. FIND THE PROBABILITY THAT THE NUMBER WHO SAYS CASHEWS ARE THEIR FAVORITE NUT (A) EXACTLY THREE, (B) MORE THAN 2, AND (C) AT MOST TWO

Answer by stanbon(75887)

Please help me solve this;46% OF ADULTS SAY CASHEWS ARE THEIR FAVORITE KIND OF NUT. YOU RANDOMLY SELECT 12 ADULTS AND ASK EACH TO NAME HIS OR HER FAVORITE NUT.
---
Binomial Problem with n = 12 and p(cashew) = 0.46
------
FIND THE PROBABILITY THAT THE NUMBER WHO SAYS CASHEWS ARE THEIR FAVORITE NUT (A) EXACTLY THREE
P(x = 3) = 12C3*0.46^3*0.54^9 = binompef(12,0.46,3) = 0.0836
==========================
(B) MORE THAN 2
P(3<= x <=12) = 1 - P(0<= x <=2) = 1-binomcdf(12,0.46,2) = 0.9637
==========================
(C) AT MOST TWO
P(0<= x <=2) = P(12,0.46,2) = bincdf(12,0.46,2) = 0.0363
===========================
Cheers,
Stan H.
===========================